OpenStudy (anonymous):

Verify the identity. PLEASE HELP!! 1- cosx/cosx/tanx =tanx/1+cosx/cosx

OpenStudy (ranga):

Click on the Equation button below the reply box and retype the identity so it is clearer.

OpenStudy (anonymous):

\[1-\cos(x)/\cos(x)/\tan(x) = \tan(x)/1+\cos(x)/\cos(x)\]

OpenStudy (ranga):

Still it not clear what is diving what. Instead of using / for divide use the fraction a/b button in the equation editor.

OpenStudy (anonymous):

\[\frac{ 1-\cos(x) }{ \cos(x) }* \frac{ 1 }{ \tan(x) } = \frac{ \tan(x) }{ 1+ \cos(x) } * \frac{ 1 }{ \cos(x) }\]

OpenStudy (ranga):

are you sure the very last cos(x) is not in the numerator?

OpenStudy (anonymous):

no its not in the numerator its like this

OpenStudy (ranga):

No. something is off on the left or the right. Left hand side is: (1 - cos(x)) / (cos(x) * tan(x)) replace tan(x) with sinx(x) / cos(x) (1 - cos(x)) / sin(x) Multiply top and bottom by (1 + cos(x)) { (1 - cos(x)) * (1 + cos(x)) } / { sin(x) * (1 + cos(x)) } (use the identity: (a + b)(a - b) = a^2 - b^2 (1 - cos^2(x)) / { sin(x) * (1 + cos(x)) } sin^2(x) / { sin(x) * (1 + cos(x)) } sin(x) / (1 + cos(x) multiply top and bottom by cos(x) sin(x) / (1 + cos(x) = { sin(x) * cos(x) } / { (1 + cos(x) * cos(x) } = tan(x) * cos(x) / (1 + cos(x)) So there is something not correct in the left or the right in problem posted.

OpenStudy (anonymous):

let me post a picture of the question

OpenStudy (ranga):

waiting on the picture ...

OpenStudy (anonymous):

http://minus.com/i/NOHBsCfjI9Wq

OpenStudy (anonymous):

number 12

OpenStudy (anonymous):

the text is very small so it doesnt show up good on camera

OpenStudy (ranga):

is one of them COT instead of COS by any chance?

OpenStudy (anonymous):

im not sure, its not very clear. would cot satisfy the equation?

OpenStudy (ranga):

there are just too many possibilities to try which one is cot. But on camera it looks smudged. but you should have a clearer view. can't you tell?

OpenStudy (anonymous):

that is precisely the problem i cant tell. frankly it looks just the same on the paper as it does on the sheet.although the first one could be cot, the second one seems like cos

OpenStudy (ranga):

where is this problem from?

OpenStudy (anonymous):

a worksheet from school

OpenStudy (ranga):

try with a magnifying glass or can you ask a friend or someone else to take a look and see what it is.

OpenStudy (ranga):

They probably copied the worksheet several times and when you make a copy of a copy of a copy it will become more and more smudged and difficult to read.

OpenStudy (anonymous):

yes thats probably it

OpenStudy (ranga):

without knowing the problem it is hard to answer.

OpenStudy (anonymous):

true say